A 10-g rifle bullets requires a speed of a 400 m/s in traversing a barrel 50 cm long. find the average acceleration and acceleration force?

1
Expert's answer
2022-05-02T13:42:20-0400

Given:

"m=0.01\:\rm kg"

"v_0=400\:\rm m/s"

"v_f=0\:\rm m/s"

"d=0.5\:\rm m"


The energy-work theorem says

"\Delta E_k=W"

"\frac{mv_0^2}{2}=Fd"

The average force

"F=\frac{mv_0^2}{2d}=\frac{0.01*400^2}{2*0.5}=1600\:\rm N"

Acceleration

"a=\frac{F}{m}=\frac{1600}{0.01}=1.6*10^5\:\rm m/s^2"